# Structured IB: Improving Information Bottleneck with Structured Feature Learning

### v1

@@ Discussion
In this section, we will demonstrate our experimental setups as well as several results.
@@
Figure 3: The accuracy, MI $I(Z,Y),I(X,Z)$ and the numbers of model parameters v.s. the number of encoders. The IB Lagrangian versions are marked by dashed lines. The upper four figures are based on MNIST while the figures in the bottom are the results of CIFAR10.
Behavior on IB Plane
Behavior of the weights
Table 1: The weights of encoders.
@@
Encoder Dropout
@@ Appendix B Implementation Details
Appendix C Supplimentary Experiments of Neural Estimators
@@
Table 3: The weights of encoders.
MNIST
K SMINE SKNIFE
2 (1.4537, 0.7509) (1.3645, 0.7112)
3 (1.6725, 0.5478, 0.7605) (1.5641, 0.6417, 0.8699)
CIFAR10
SMINE SKNIFE
2 (2.0496, 0.4243) (1.8342, 0.2867)
3 (1.1864, -0.5831, 1.3554) (2.1334, 0.7671, 0.4313)
@@
Input: Training set $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$, encoders $E,\{E_{i}\}_{i=1}^{K}$, decoder $D$, discriminator $d$, weights $\{w_{i}\}_{i=0}^{K}$, maximal epoch $epoch$ and $\beta$.
@@
1:  Let $e=0$;
2:  for $e\leq epoch$ do
3:     for each random batch $\{x_{j},y_{j}\}_{j\in\mathcal{B}}$ do
4:        train $E$ and $D$ by $\min_{E,D}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-I(E(x_{j}),y_{j})+\beta I(x_{j},E(x_{j}));$
@@
6:  end for
7:  Let $e=0$;
8:  for $e\leq epoch$ do
9:     for $i\in[1,K]$ do
10:        for each random batch $\{x_{j},y_{j}\}_{j\in\mathcal{B}}$ do
11:           Obtaining samples $\{z_{ij}=E_{i}(x_{j}),z_{j}=E(x_{j}),z_{lj}=E_{l}(x_{j})\}_{j\in\mathcal{B},l\in[1,i-1]};$
12:           Training $E_{i}$ by $\min_{E_{i}}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-I(z_{ij},y_{j})+\beta I(x_{j},z_{ij})-\log d(z_{ij},z+\sum_{l\in[1,i-1]}z_{lj});$
13:           Random shuffling on batch indices $\mathcal{B}$, obtaining $\{z_{i\pi(j)},z_{\pi(j)}+\sum_{l\in[1,i-1]}z_{l\pi(j)}\}_{j\in\mathcal{B},l\in[1,i-1]}$ where $\pi()$ is the shuffling projection;
14:           Training $d$ by $\min_{d}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-\log(1-d(z_{ij},z+\sum_{l\in[1,i-1]}z_{lj}))-\log d(z_{i\pi(j)},z_{\pi(j)}+\sum_{l\in[1,i-1]}z_{l\pi(j)});$
@@
16:     end for
17:  end for
18:  for $e\leq epoch$ do
19:     for each random batch $\{x_{j},y_{j}\}_{j\in\mathcal{B}}$ do
20:        Calculate $\hat{z}_{j}=w_{0}E(x_{j})+\sum_{i=1}^{K}w_{i}E_{i}(x_{j});$
21:        Training weights $\{w_{i}\}_{i=0}^{K}.$ by $\min_{\{w_{i}\}_{i=0}^{K}}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-I(\hat{z}_{j},y_{j})+\beta I(x_{j},\hat{z}_{j});$
22:     end for
23:  end for
24:  return $E,\{E_{i}\}_{i=1}^{K}$, $D$ and $\{w_{i}\}_{i=0}^{K}$.

### v2

@@ Discussion
In this section, we will demonstrate our experimental setups and several results. Specifically, we first evaluate and analyze the performance of the proposed SIB in comparison with its corresponding IB algorithm. Next, we examine the behavior of weights in the SIB, followed by an analysis of both SIB and corresponding IB methods on the IB plane. Due to space limitations, experiments on encoder dropout, which show the significance of SIB weights, are provided in the Appendix.
@@
Table 1: The weights of encoders.
Figure 3: The accuracy, MI $I(Z,Y),I(X,Z)$ and the numbers of model parameters (Num. Params., in Millions) v.s. the number of encoders. The IB Lagrangian versions are marked by dashed lines. The left four figures are based on MNIST while the figures in the right are the results of CIFAR10.
Behavior of Weights
Behavior on IB Plane
@@ Appendix B Implementation Details
Appendix C Encoder Dropout
Appendix D Supplimentary Experiments of Neural Estimators
@@
MNIST
K SMINE SKNIFE
2 (1.453, 0.751) (1.365, 0.711)
3 (1.673, 0.548, 0.761) (1.564, 0.642, 0.870)
CIFAR10
SMINE SKNIFE
2 (2.050, 0.424) (1.834, 0.287)
3 (1.186, -0.583, 1.355) (2.133, 0.767, 0.431)
Table 3: The weights of encoders.
@@
Input: Training set $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$, encoders $E,\{E_{i}\}_{i=1}^{K}$, decoder $D$, discriminator $d$, weights $\{w_{i}\}_{i=0}^{K}$, $epoch$ and $\beta$.
@@
1:  for $e=1\text{ to }epoch$ do
2:     for each random batch $\{x_{j},y_{j}\}_{j\in\mathcal{B}}$ do
3:        train $E$ and $D$ by $\min_{E,D}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-I(E(x_{j}),y_{j})+\beta I(x_{j},E(x_{j}));$
4:     end for
@@
6:  for $e=1\text{ to }epoch$ do
7:     for $i\in[1,K]$ do
8:        for each random batch $\{x_{j},y_{j}\}_{j\in\mathcal{B}}$ do
9:           Obtaining samples $\{z_{ij}=E_{i}(x_{j}),z_{j}=E(x_{j}),z_{lj}=E_{l}(x_{j})\}_{j\in\mathcal{B},l\in[1,i-1]};$
10:           Training $E_{i}$ by $\min_{E_{i}}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-I(z_{ij},y_{j})+\beta I(x_{j},z_{ij})-\log d(z_{ij},z+\sum_{l\in[1,i-1]}z_{lj});$
11:           Shuffling on batch indices $\mathcal{B}$, obtaining $\{z_{ij},z_{\pi(j)}+\sum_{l\in[1,i-1]}z_{l\pi(j)}\}_{j\in\mathcal{B}}$ where $\pi()$ is the shuffling projection;
12:           Training $d$ by $\min_{d}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-\log(1-d(z_{ij},z+\sum_{l\in[1,i-1]}z_{lj}))-\log d(z_{i\pi(j)},z_{\pi(j)}+\sum_{l\in[1,i-1]}z_{l\pi(j)});$
13:        end for
14:     end for
@@
16:  for $e=1\text{ to }epoch$ do
17:     for each random batch $\{x_{j},y_{j}\}_{j\in\mathcal{B}}$ do
18:        Calculate $\hat{z}_{j}=w_{0}E(x_{j})+\sum_{i=1}^{K}w_{i}E_{i}(x_{j});$
19:        Training weights $\{w_{i}\}_{i=0}^{K}.$ by $\min_{\{w_{i}\}_{i=0}^{K}}\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}-I(\hat{z}_{j},y_{j})+\beta I(x_{j},\hat{z}_{j});$
20:     end for
21:  end for
22:  return $E,\{E_{i}\}_{i=1}^{K}$, $D$ and $\{w_{i}\}_{i=0}^{K}$.